∫(tan^-1 x)/(1+x² )dx from 0 to 1


1
Expert's answer
2021-03-01T06:40:58-0500

"\displaystyle\int_0^1\frac{\tan^{-1}x}{1+x^2}\,dx \\
= \int_0^1\tan^{-1}x\,\frac d{dx}(\tan^{-1}x)\,dx \\
= \frac12(\tan^{-1}x)^2\Big|_0^1 \\
= \frac12[(\tan^{-1}1)^2 - \tan^{-1}0)^2] \\
= \frac12[(\frac \pi4)^2 - 0] \\
= \frac {\pi^2}{32}"